# Controllable Human-centric Keyframe Interpolation with Generative Prior

### v1

@@ 3.2 VDMs as Base Interpolation Engine
To supply generative priors for human-centric keyframe interpolation, we adapt pre-trained Video Diffusion Models (VDMs) under the latent diffusion framework [33]. VDMs perform the diffusion process in a VAE-encoded latent space, conditioning video synthesis on input frames. Our main experiments investigate Wan2.1 [37] as the base model for keyframe interpolation.
Wan2.1 consists of a scaled-up DiT-based denoiser [29] and a causal 3D-VAE that performs spatiotemporal compression. It employs the flow matching strategy [22, 9] for the diffusion process. It formulates the forward diffusion process as a linear interpolation between the clean video latent z and the noise $\bm{\epsilon}$, which adds the noise by: $\textbf{z}_{t}=(1-t)\textbf{z}+t\bm{\epsilon}$. In the backward process, the denoiser $f_{\theta}$ iteratively refines $\mathbf{z}_{t}$ conditioned on the first frame $I_{0}$. The training objective is formulated as:
$\mathcal{L}=\mathbb{E}_{\textbf{z},t,I_{0},\bm{\epsilon}\sim\mathcal{N}(0,I)}[||f_{\theta}(\textbf{z}_{t},I_{0},t)-\textbf{y}||_{2}^{2}],$ (6)
where the target objective y is $\frac{d\textbf{z}_{t}}{dt}=\bm{\epsilon}-\textbf{z}$. Since the latent space is unevenly compressed in time, noise fusion strategies [10, 52] that combine temporally forward and reverse denoising paths are ineffective. Instead, we adapt Wan2.1 for keyframe interpolation in one unified denoising process. Specifically, we condition Wan2.1 on both endpoint frames $I_{0},I_{N}$, and apply LoRA tuning to its input patch embeddings. Additional details are provided in the supplementary material.
@@ 4 The CHKI-Video Dataset
Existing interpolation datasets [43, 7, 5, 34, 36], which are designed for interpolating temporally adjacent frames, are not suitable for the CHKI task. To address this gap, we built the CHKI-Video dataset in three consecutive stages. More details can be found in the supplementary material.
Stage 1: Dataset Collection. To create a dataset with challenging and diverse human motion, we curate video clips from SportSlomo [5] and Pexels [^2] https://www.pexels.com/videos/. While SportsSlomo provides challenging human-centric videos, its exclusive focus on sports limits the diversity of activities. To enhance diversity, we compile a list of keywords, spanning everyday activities to high-intensity actions. We use these keywords to retrieve videos from Pexels, complementing the sports videos. The curated videos are then downsampled to eliminate frame redundancy unnecessary for keyframe interpolation.
Stage 2: Pre-annotation Processing. We first perform general filtering for the low-quality videos according to brightness changes and assessed scores [40]. Then, we use Grounding-DINO [23] and SAM2 [32] to detect, segment, and track human instances in each video. We discard any video of more than three people or is shorter than 20 consecutive frames to guarantee a sufficient temporal span for keyframe interpolation. After automated processing, we manually review and filter detections in complex sports scenarios.
Stage 3: Human-centric Annotation. Building on the accurate human detections obtained in Stage 2, we annotate each clip for precise human-centric information. First, we employ Sapiens [18] to extract 2D human keypoints and perform whole-body detection to filter out clips with incomplete human figures. This ensures our dataset remains strictly human-centric. Finally, we apply SMPLer-[3], leveraging its high re-projection accuracy to fit detailed SMPL-X models for human images and produce reliable 3D body parameters for each frame.
As a result, CHKI-Video comprises 2,614 video clips of over 180K frames, carefully annotated with human bounding boxes, masks, 2D keypoints and 3D parametric models. To prepare the train and test split, we follow the original division for SportsSlomo [5] videos and distribute the Pexels videos according to their keyword frequencies to maintain balanced coverage of all motion categories.
@@ 5 Experiments
We present quantitative and qualitative results in Sec. 5.1 to validate the effectiveness of our 3D control strategy in PoseFuse3D. We compare our interpolation performance against state-of-the-art methods in Sec. 5.2. Finally, Sec. 5.3 provides a detailed ablation study to justify our model design.
@@
Evaluation Protocols. We evaluate methods on CHKI-Video’s test set using the ground-truth annotations to drive controllable interpolation. With motion ambiguity limited by the ground truth controls, we adopt the standard interpolation metrics of PSNR and LPIPS computed on the whole images for evaluation. To quantify performance specifically on human regions, we further leverage the annotated human boxes and masks, yielding PSNRbbox, LPIPSbbox, PSNRmask, and LPIPSmask metrics. Notably, we apply binary dilation to expand and smooth the masks before computing the metrics, preventing potential artifacts along mask boundaries.
@@
FCVG SVD N.A. 20.42 13.28 12.14 0.2100 0.0899 0.0606
PoseFuse3D SVD VE-SVD 20.96 13.79 12.70 0.1975 0.0835 0.0553
PoseFuse3D Wan2.1-I2V VE 21.91 14.36 13.10 0.1400 0.0682 0.0484
PoseFuse3D Wan2.1-I2V VE+DN 22.07 14.35 13.03 0.1363 0.0667 0.0473
PoseFuse3D Wan2.1-I2V VE+SE (Ours) 22.14 14.53 13.24 0.1330 0.0653 0.0464
@@
Setup. We evaluate the effectiveness of the 3D-informed control design in PoseFuse3D via comparisons among 3D control strategies, including ‘VE’, ‘VE+DN’ and ‘VE+SE’. Specifically, VE refers to the visual encoding on the SMPL-X colored surface rendering, removing the SMPL-X encoder from the full design of our control model. VE+DN extends VE by incorporating depth and normal renderings, using two additional encoders identical to the one used in VE. VE+SE represents our proposed strategy in PoseFuse3D that directly encodes 3D information via the SMPL-X encoder and integrates it with VE. Experiments for these strategies use Wan2.1-I2V [37] as the interpolation backbone. For efficiency, they are trained for 40K iterations.
To assess the necessity of 3D information, we compare against FCVG [52], which is conditioned solely on 2D signals. For fair comparison, we create a variant of VE by replacing the backbone with SVD [2]. We adapt SVD for interpolation with the temporal forward and reverse denoising path fusion strategy used in FCVG. This variant is referred to as VE-SVD for ease of analysis.
@@ 5.2 Benchmark Results
GIMM-VFI [11] 20.29 11.59 10.16 0.1954 0.1187 0.0860 0.9146
GI [39] 15.81 8.27 7.26 0.3364 0.1672 0.1146 0.8954
Wan2.1-KI (Ours) 19.53 11.19 9.85 0.2081 0.1208 0.0868 0.9180
FCVG [52] 20.42 13.28 12.14 0.2100 0.0899 0.0606 0.9187
PoseFuse3D-KI (Ours) 22.27 14.72 13.47 0.1304 0.0636 0.0450 0.9189
@@
Conv-Enc (2D) N.A. 20.29 12.17 11.33 0.1990 0.1096 0.0836
Dual Conv-Enc Sum 20.50 12.43 11.59 0.1953 0.1059 0.0806
Dual Conv-Enc Non-TSA 20.30 12.19 11.35 0.1956 0.1053 0.0790
Dual Conv-Enc Non-WP 20.46 12.42 11.60 0.1938 0.1038 0.0775
Dual Conv-Enc Full 20.55 12.53 11.71 0.1927 0.1031 0.0773
Figure 4: Qualitative Comparisons with State-of-The-Art Methods.
Setup. We compare PoseFuse3D-KI against several advanced interpolation methods on our CHKI-Video dataset. The main comparison is with FCVG [52], which also enables intermediate control during interpolation. For broader coverage, we also include the keyframe interpolation method GI [39] and the traditional video frame interpolation method GIMM-VFI [11]. We also include Wan2.1-KI, an adaptation of Wan2.1 [37] for keyframe interpolation, following the strategy in Sec. 3.2. Since not all baselines support ground truth controls, we additionally compute the Human Anatomy (HA) [49] score to assess the quality of human synthesis and support fair comparison.
Results. We report quantitative results of PoseFuse3D-KI on the CHKI-Video benchmark in Table 2. Our method delivers state-of-the-art performance on human-centric keyframe interpolation. On whole-image metrics, it boosts PSNR by 1.85 dB and lowers LPIPS by 0.0796 in comparison with the state-of-the-art method FCVG [52]. Crucially, it also outperforms on human-centric metrics compared with other methods, achieving a PSNRbbox of 14.72 dB, an LPIPSmask of 0.045, and an HA score of 0.9189. This indicates that our method produces plausible, high-fidelity human interpolations that closely follow the ground-truth dynamics, demonstrating the effectiveness of our method.
Visualizations. For qualitative evaluation, we qualitatively compare PoseFuse3D-KI with other advanced methods, as illustrated in Figure 4. Our approach delivers more accurate human interpolations, faithfully following real-world motions and preserving body shape. For example, in the $2^{nd}$ ‘Fencing’ case and $4^{th}$ ‘Stunt Bike’ case, only PoseFuse3D-KI correctly interpolates leg and arm movements while maintaining consistent shapes. Moreover, our method naturally handles the occluded human motion well with 3D-informed control. We observe that our method interpolates correct spatial positions for the occluded legs ($1^{st}$ case) and arms ($3^{nd}$ case), achieving significant improvements over FCVG. Furthermore, although the control-free keyframe interpolation methods GI and Wan2.1-KI occasionally generate undistorted humans, they often generate implausible motion that violates real-world dynamics, as observed in $1^{st}$, $2^{nd}$, and $4^{th}$ cases.
5.3 Ablation Study
Fusion Module. In PoseFuse3D, condition features are fused through a carefully designed fusion module. To validate its effectiveness, we replace it with a simple summation operation [51], denoted as ‘Sum’ in Table 3. This change leads to a significant performance drop, particularly in perceptual quality, with an increase of 0.0033 in LPIPSmask. These results demonstrate the fusion module’s contribution to providing informative control for high-quality interpolation.

### v2

@@ 3.2 VDMs as Base Interpolation Engine
To supply generative priors for human‑centric keyframe interpolation, we adapt pre‑trained Video Diffusion Models (VDMs) under the latent diffusion framework [rombach2022high]. VDMs perform the diffusion process in a VAE‑encoded latent space, conditioning video synthesis on input frames. Our main experiments investigate Wan2.1 [wang2025wan] as the base model for keyframe interpolation.
Wan2.1 consists of a scaled‑up DiT‑based denoiser [peebles2023dit] and a causal 3D-VAE that performs spatiotemporal compression. It employs the flow matching strategy [lipmanflowmatching, esser2024rectifiedflow] for the diffusion process. It formulates the forward diffusion process as a linear interpolation between the clean video latent z and the noise $\bm{\epsilon}$, which adds the noise by: $\textbf{z}_{t}=(1-t)\textbf{z}+t\bm{\epsilon}$. In the backward process, the denoiser $f_{\theta}$ iteratively refines $\mathbf{z}_{t}$ conditioned on the first frame $I_{0}$. The training objective is formulated as:
[…]
where the target objective y is $\frac{d\textbf{z}_{t}}{dt}=\bm{\epsilon}-\textbf{z}$. Since the latent space is unevenly compressed in time, noise fusion strategies [feng2024trf, zhu2024generative] that combine temporally forward and reverse denoising paths are ineffective. Instead, we adapt Wan2.1 for keyframe interpolation in one unified denoising process. Specifically, we condition Wan2.1 on both endpoint frames $I_{0},I_{N}$, and apply LoRA tuning to its input patch embeddings. Additional details are provided in the supplementary material.
@@ 4 The CHKI-Video Dataset
Existing interpolation datasets [xue2019vimeo90k, choi2020snufilm, chen2024sportsslomo, sim2021xvfi, stergiou2024lavib], which are designed for interpolating temporally adjacent frames, are not suitable for the CHKI task. To address this gap, we built the CHKI‑Video dataset in three consecutive stages. More details can be found in the supplementary material.
Stage 1: Dataset Collection. To create a dataset with challenging and diverse human motion, we curate video clips from SportSlomo [chen2024sportsslomo] and Pexels [pexels]. While SportsSlomo provides challenging human-centric videos, its exclusive focus on sports limits the diversity of activities. To enhance diversity versus the reality [yu2023dataset], we compile a list of keywords, spanning everyday activities to high-intensity actions. We use these keywords to retrieve videos from Pexels, complementing the sports videos. The curated videos are then downsampled to eliminate frame redundancy unnecessary for keyframe interpolation.
Stage 2: Pre-annotation Processing. We first perform general filtering for the low-quality videos according to brightness changes and assessed scores [wu2023dover]. Then, we use Grounding-DINO [liu2024groundingdino] and SAM2 [ravi2024sam2] to detect, segment, and track human instances in each video. We discard any video of more than three people or that is shorter than 20 consecutive frames to ensure a sufficient temporal span for keyframe interpolation. After automated processing, we manually review and filter detections in complex sports scenarios.
Stage 3: Human-centric Annotation. Building on the accurate human detections obtained in Stage 2, we annotate each clip for precise human-centric information. First, we employ Sapiens [khirodkar2024sapiens] to extract 2D human keypoints and perform whole‑body detection to filter out clips with incomplete human figures. This ensures our dataset remains strictly human-centric. Finally, we apply SMPLer‑X [cai2023smpler], leveraging its high re‑projection accuracy to fit detailed SMPL‑X models for human images and produce reliable 3D body parameters for each frame.
As a result, CHKI-Video comprises 2,614 video clips of over 180K frames, carefully annotated with human bounding boxes, masks, 2D keypoints and 3D parametric models. To prepare the train and test split, we follow the original division for SportsSlomo [chen2024sportsslomo] videos and distribute the Pexels videos according to their keyword frequencies to maintain balanced coverage of all motion categories.
@@ 5 Experiments
We present quantitative and qualitative results in Sec. 5.1 to validate the effectiveness of our 3D control strategy in PoseFuse3D. We compare our interpolation performance against state-of-the-art methods in Sec. 5.2 and analyze the scalability across temporal gaps in Sec. 5.3. We further assess the in-the-wild interpolation capability in Sec. 5.4, where ground-truth control signals are not available. Finally, Sec. 5.5 provides a detailed ablation study to justify our model design.
@@
Evaluation Protocols. Unless otherwise noted, all methods are evaluated on the CHKI-Video test set using ground-truth annotations for controllable interpolation. With motion ambiguity limited by the ground truth controls, we adopt the standard interpolation metrics of PSNR and LPIPS computed on the whole images for evaluation. To quantify performance specifically on human regions, we further leverage the annotated human boxes and masks, yielding PSNRbbox, LPIPSbbox, PSNRmask, and LPIPSmask metrics. Notably, we apply binary dilation to expand and smooth the masks before computing the metrics, preventing potential artifacts along mask boundaries.
@@
FCVG SVD N.A. 20.42 18.05 16.91 0.2100 0.0899 0.0606
PoseFuse3D SVD VE-SVD 20.96 18.56 17.47 0.1975 0.0835 0.0553
PoseFuse3D Wan2.1-I2V VE 21.91 19.13 17.87 0.1400 0.0682 0.0484
PoseFuse3D Wan2.1-I2V VE+DN 22.07 19.12 17.80 0.1363 0.0667 0.0473
PoseFuse3D Wan2.1-I2V VE+SE (Ours) 22.14 19.30 18.01 0.1330 0.0653 0.0464
@@
Setup. We evaluate the effectiveness of the 3D-informed control design in PoseFuse3D via comparisons among 3D control strategies, including ‘VE’, ‘VE+DN’ and ‘VE+SE’. Specifically, VE refers to the visual encoding on the SMPL-X colored surface rendering, removing the SMPL-X encoder from the full design of our control model. VE+DN extends VE by incorporating depth and normal renderings, using two additional encoders identical to the one used in VE. VE+SE represents our proposed strategy in PoseFuse3D that directly encodes 3D information via the SMPL-X encoder and integrates it with VE. Experiments for these strategies use Wan2.1-I2V [wang2025wan] as the interpolation backbone. For efficiency, they are trained for 40K iterations.
To assess the necessity of 3D information, we compare against FCVG [zhu2024generative], which is conditioned solely on 2D signals. For fair comparison, we create a variant of VE by replacing the backbone with SVD [blattmann2023svd]. We adapt SVD for interpolation with the temporal forward and reverse denoising path fusion strategy used in FCVG. This variant is referred to as VE-SVD for ease of analysis.
@@ 5.2 Benchmark Results
GIMM-VFI [guo2025generalizable] 20.29 16.36 14.93 0.1954 0.1187 0.0860 0.9146
GI [wang2024generative] 15.81 13.04 12.03 0.3364 0.1672 0.1146 0.8954
Wan2.1-KI (Ours) 19.53 15.96 14.62 0.2081 0.1208 0.0868 0.9180
FCVG [zhu2024generative] 20.42 18.05 16.91 0.2100 0.0899 0.0606 0.9187
PoseFuse3D-KI (Ours) 22.27 19.49 18.24 0.1304 0.0636 0.0450 0.9189
Figure 4: Qualitative Comparisons with State-of-The-Art Methods.
Setup. We compare PoseFuse3D-KI against several advanced interpolation methods on our CHKI-Video dataset. The main comparison is with FCVG [zhu2024generative], which also enables intermediate control during interpolation. For broader coverage, we also include the keyframe interpolation method GI [wang2024generative] and the traditional video frame interpolation method GIMM-VFI [guo2025generalizable]. We also include Wan2.1-KI, an adaptation of Wan2.1 [wang2025wan] for keyframe interpolation, following the strategy in Sec. 3.2. Since not all baselines support ground truth controls, we additionally compute the Human Anatomy (HA) [zheng2025vbench] score to assess the quality of human synthesis and support fair comparison.
Results. We report quantitative results of PoseFuse3D‑KI on the CHKI‑Video benchmark in Table 2. Our method delivers state‑of‑the‑art performance on human‑centric keyframe interpolation. On whole‑image metrics, it boosts PSNR by 1.85 dB and lowers LPIPS by 0.0796 in comparison with the state-of-the-art method FCVG [zhu2024generative]. Crucially, it also outperforms on human-centric metrics compared with other methods, achieving a PSNRbbox of 19.49 dB, an LPIPSmask of 0.045, and an HA score of 0.9189. This indicates that our method produces plausible, high‑fidelity human interpolations that closely follow the ground‑truth dynamics, demonstrating the effectiveness of our method.
Visualizations. For qualitative evaluation, we qualitatively compare PoseFuse3D-KI with other advanced methods, as illustrated in Figure 4. Our approach delivers more accurate human interpolations, faithfully following real-world motions and preserving body shape. For example, in the $2^{nd}$ ‘Fencing’ case and $4^{th}$ ‘Stunt Bike’ case, only PoseFuse3D‑KI correctly interpolates leg and arm movements while maintaining consistent shapes. Moreover, our method naturally handles the occluded human motion well with 3D-informed control. We observe that our method interpolates correct spatial positions for the occluded legs ($1^{st}$ case) and arms ($3^{nd}$ case), achieving significant improvements over FCVG. Furthermore, although the control-free keyframe interpolation methods GI and Wan2.1-KI occasionally generate undistorted humans, they often generate implausible motion that violates real‑world dynamics, as observed in $1^{st}$, $2^{nd}$, and $4^{th}$ cases.
5.3 Scalability Across Temporal Gaps
Table 3: Results across Temporal Gaps.
Temporal Gap Metrics
[…]
24-frame 23.86 21.64 20.70 0.1144 0.0508 0.0332
48-frame 21.46 19.48 18.57 0.1566 0.0644 0.0432
64-frame 20.37 18.34 17.45 0.1882 0.0755 0.0514
96-frame 19.44 17.50 16.66 0.2208 0.0836 0.0577
Setup. We evaluate the scalability of PoseFuse3D-KI in interpolating frames across different temporal gaps. The temporal gap is defined as the number of frames between the input keyframes, ranging from 24 to 96 frames. Evaluations are conducted on the CHKI-Video test set using PSNR and LPIPS as quantitative metrics.
Results. We provide quantitative results of PoseFuse3D-KI across different temporal gaps in Table 3. As a common trend observed in most interpolation methods, the interpolation performance gradually declines as the interval between keyframes increases. Nevertheless, PoseFuse3D-KI maintains strong performance over a wide range of temporal settings. Notably, it achieves a PSNR of 23.86 dB and an LPIPS of 0.1144 at 24 frames, and a PSNR of 21.46 dB and an LPIPS of 0.1566 at 48 frames. Even at a large gap of 96 frames, where the task becomes substantially more challenging, PoseFuse3D-KI remains competitive, achieving a PSNR of 19.44 dB and an LPIPS of 0.2208. These results demonstrate the robustness of PoseFuse3D-KI under long-range controllable interpolation.
5.4 In-the-wild Interpolation Results
Our PoseFuse3D-KI framework can be readily applied to interpolate in-the-wild human-centric keyframes without ground-truth control signals.
Setup. We assess the in-the-wild interpolation capability of PoseFuse3D-KI on the CHKI-Video dataset by discarding the ground-truth annotations. The comparison is conducted against the state-of-the-art method FCVG [zhu2024generative]. Since ground-truth controls are unavailable in this setting, we employ a simple strategy that linearly interpolates human body joints to generate intermediate control signals. Details of this pipeline are provided in the supplementary material. For performance measurement, we compute the same metrics as described in Sec. 5.2.
Results. We report quantitative results of PoseFuse3D-KI under the in-the-wild interpolation setting in Table 4. Our method achieves state-of-the-art performance and delivers significant improvements over FCVG [zhu2024generative]. On whole-image metrics, it boosts PSNR by 4.5% and reduces LPIPS by 22.1% compared with FCVG. More importantly, PoseFuse3D-KI attains the best results on perceptual human-centric metrics, with an LPIPSmask of 0.0859 and an HA score of 0.9289. These results demonstrate that PoseFuse3D-KI effectively handles in-the-wild interpolation.
5.5 Ablation Study
Fusion Module. In PoseFuse3D, condition features are fused through a carefully designed fusion module. To validate its effectiveness, we replace it with a simple summation operation [zhu2024champ], denoted as ‘Sum’ in Table 5. This change leads to a significant performance drop, particularly in perceptual quality, with an increase of 0.0033 in LPIPSmask. These results demonstrate the fusion module’s contribution to providing informative control for high-quality interpolation.
Table 4: Comparisons with FCVG on In-the-wild Interpolation.
Methods Metrics
[…]
FCVG [zhu2024generative] 18.47 15.31 14.08 0.2607 0.1321 0.0921 0.9284
PoseFuse3D-KI (Ours) 19.30 15.75 14.46 0.2031 0.1194 0.0859 0.9289
@@ 5.5 Ablation Study
Conv-Enc (2D) N.A. 20.29 16.94 16.10 0.1990 0.1096 0.0836
Dual Conv-Enc Sum 20.50 17.20 16.36 0.1953 0.1059 0.0806
Dual Conv-Enc Non-TSA 20.30 16.96 16.12 0.1956 0.1053 0.0790
Dual Conv-Enc Non-WP 20.46 17.19 16.37 0.1938 0.1038 0.0775
Dual Conv-Enc Full 20.55 17.30 16.48 0.1927 0.1031 0.0773